# bibliography.bib

### v2

@@ 1 Introduction
In a lot of machine learning problems we want to embed our data into an Euclidean space $\mathbb{R}^{m}$ using a symmetry-invariant embedding $\Psi$ and utilize $\mathbb{R}^{m}$ as our feature space. This embedding $\Psi$ should also separate data orbits and satisfy certain stability conditions to ensure that small perturbations of the input do not significantly impact the predictions. We worked in the orbit separation problem in [balan2023inj]. In this paper we focus on the stability problem.
This problem is an instance of invariant machine learning. [mixon2022injectivity, aslan2023group, dym2022low, bronstein, maron2018invariant, DUFRESNE20091979, cahill2023bilipschitz].
The most common group actions in invariant machine learning are permutations [sannai2020universal, cahill22, balan2022permutation] reflections [mixon2022max] and translations [Cahill19]. A related, by somewhat different problem is the case of equivariant embeddings [lipman2022, maron2018invariant, villar2021scalars, villar2022dimensionless, blum2022equivariant].
Our work extends and unifies two prior approaches: the max filter embedding introduced in [cahill22, mixon2022injectivity], and the permutation invariant representation introduced in [balan2022permutation].
The construction of permutation invariant embeddings is closely connected to the phase retrieval problem [balan2004signal, bcmn] that has a large body of results. For instance, [balan16] provides exact estimates for both the upper and lower Lipschitz bounds, addressing both the real and complex cases of the phase retrieval problem. In [Balan2023Rel] we establish an isometric identification of the real phase retrieval problem to $S_{2}$-invariant representations.
A completely different approach is considered in [eriksson2018quantitative]. There it is proved that for any discrete group $G$ of isometries acting on $\mathbb{R}^{d}$, it is possible to construct a bi-Lipschitz map from $\mathbb{R}^{d}/G$ to $\mathbb{R}^{N}$, where the distortion depends solely on $d$.
Since the first draft of this paper was placed on arxiv, the authors of [qaddura2024stable] extended the construction of this paper and proved that for certain compact groups, given enough generic templates, the coorbit filter bank (as defined there) is injective and bi-Lipschitz.
@@ 1.1 Notation
In this paper we use the same notation as in [balan2023inj]. Let $({\mathscr{V}},{\langle\cdot,\cdot\rangle})$ be a $d$-dimensional real vector space with scalar product, and $d\geq 2$. Assume $(G,\cdot)$ is a finite group of order $|G|=N$ acting unitarily on ${\mathscr{V}}$. For every $g\in G$, we denote by $U_{g}x$ the group action on vector $x\in{\mathscr{V}}$. Let $\hat{{\mathscr{V}}}={\mathscr{V}}/\sim$ denote the quotient space with respect to this group action, $x\sim y$ if and only if $y=U_{g}x$ for some $g\in G$. We denote by $[x]$ the orbit of vector $x$, i.e. $[x]=\{U_{g}x:g\in G\}$. The natural metric, $\operatorname{{\mathbf{d}}}:\hat{{\mathscr{V}}}\times\hat{{\mathscr{V}}}\to\mathbb{R}$, is defined by
@@
For an integer $p\in\mathbb{N}$, we denote $[p]=\{1,2,\ldots,p\}$. For a set $S$, $|S|$ denotes its cardinal. Fix a $p$-tuple of vectors ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$. For any $i\in[p]$ and $j\in[N]$ we define the operator $\Phi_{w_{i},j}:{\mathscr{V}}\to\mathbb{R}$ so that $\Phi_{w_{i},j}(x)$ is the $j$-th coordinate of the sorted vector $\operatorname{\downarrow}({\langle U_{g}w_{i},x\rangle})_{g\in G}$. Fix a set $S\subset[N]\times[p]$ such that $|S|=m$, and for $i\in[p]$, let $S_{i}=\{k\in[N]:(k,i)\in S\}$ (the $i^{th}$ column of $S$). Denote by $m_{i}$ the cardinal of the set $S_{i}$, $m_{i}=|S_{i}|$. Thus $m=\sum_{i=1}^{p}m_{i}$.
@@
$[y]$$[x]$$\hat{{\mathscr{V}}}$$\mathbb{R}^{m}$$\mathbb{R}^{q}$$\Phi_{{\textbf{w}},S}$$\Phi_{{\textbf{w}},S}$$\ell$$\ell$$\Phi_{{\textbf{w}},S}([x])$$\Phi_{{\textbf{w}},S}([y])$$\ell\circ\Phi_{{\textbf{w}},S}([y])$$\ell\circ\Phi_{{\textbf{w}},S}([x])$
@@
Informally, our main results, 1.4, 1.6, 1.7 and 1.8, state that: (1) "injectivity" implies "(bi-Lipschitz) stability", (2) stable bi-Lipschitz embedding can be achieved into an Euclidean space of dimension at most twice the dimension of the input space; and (3) any continuous or even Lipschitz map factors through $\Phi_{{\textbf{w}},S}$.
@@ 1.2 Main results
Assume that, for fixed $S\in[N]\times[p]$ and ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$, the map $\hat{\Phi}_{{\textbf{w}},S}:{\mathscr{V}}\to\mathbb{R}^{m}$ is bi-Lipschitz on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/G$. Let $L$ denote a locally convex topological vector space. For every continuous map $F:\hat{{\mathscr{V}}}\to L$ there exists a continuous $T:\mathbb{R}^{m}\to L$ such that $F=T\circ\Phi_{{\textbf{w}},S}$. Furthermore $T(\mathbb{R}^{m})$ is included in the convex hull of $F(\hat{{\mathscr{V}}})$.
@@ 2 Stability of Embedding
Suppose that for ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$ the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective. In this case, we claim the map $\hat{\Phi}_{{\textbf{w}},S}$ is also bi-Lipschitz, (1.4). An estimate of the upper Lipschitz constant $b$ is given in 2.1.
1.5 follows from 1.4.
The proof of 1.4 is contained in the next two subsections.
@@ 2.1 Upper Lipschitz bound
[…]
@@ 2.2.1 Geometric Analysis of Coorbits
$\Delta^{i,j}(x)=\begin{cases}\min_{g\notin L^{i,j}(x)}(|{\langle U_{g}w_{i},x\rangle}-\Phi_{i,j}(x)|)\frac{1}{\lVert w_{i}\rVert},\leavevmode\nobreak\ \text{ if }L^{i,j}(x)\not=G\\ \frac{\lVert x\rVert}{\lVert w_{i}\rVert},\leavevmode\nobreak\ \text{ if }L^{i,j}(x)=G.\\ \end{cases}$ (5)
@@
then there exist at least $N-j+1$, distinct elements of group $G$, say $\{h_{1},\dots h_{j}\}$ such that ${\langle U_{h_{k}}w_{i},x\rangle}<{\langle U_{g}w_{i},x\rangle}\leavevmode\nobreak\ \forall k\in[N-j+1]$, which is also a contradiction.
@@
But the set $\cup_{k\leq j}L^{i,k}(x)$ contains at least $j$ elements (since each $L^{i,j}(x)$ is non-empty) and so we derived a contradiction with 2.2(a) Equation 6.
@@
Assume that exist $g_{1}\in L^{i,j}(x+c_{2}y)$ with $g_{1}\notin L^{i,j}(x+c_{1}y)$. Without loss of generality assume that ${\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}<\Phi_{i,j}(x+c_{1}y)$. Let $q>j$ be the smallest integer such that $g_{1}\in L^{i,q}(x+c_{1}y)$. Then $\Phi_{i,q}(x+c_{1}y)={\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}<\Phi_{i,j}(x+c_{1}y)$. By 2.2 (b)(ii),
@@
and $g_{1}\not\in\cup_{r\leq j}L^{i,r}(x+c_{1}y)$. On the other hand, from 2.2 (a), Equation 6,
@@
1. Note that by norm conditions on $z_{r}$, 2.3 and 2.4, if $g_{1},g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$ then ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}\leavevmode\nobreak\ \forall l\in[k]$. Starting with $l=1$ and proceeding recursively we get ${\langle U_{g_{1}}w_{i},z_{r}\rangle}={\langle U_{g_{2}}w_{i},z_{r}\rangle}\leavevmode\nobreak\ \forall r\in[k]$. Therefore ${\langle U_{g_{1}}w_{i},a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},a_{r}z_{r}\rangle}\leavevmode\nobreak\ \forall r\in[k]$ and so ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}.$
@@
$\subset$". From 2.3 we have that $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$.
$\supset$". Let $g_{1}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$. Take any $g_{2}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. By 2.3, $g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$, but from part 1 of this Lemma,
@@
Case 1. Suppose $L^{i,j}(\sum_{r=1}^{k}z_{r})=G$. From 2.3 we know that
@@ 5. We prove the equality between complements:
First notice that by 2.3 we have that $L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$ and $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. From 2.6 part (2) we have that $L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. Take $g\in L^{i,j}(\sum_{r=1}^{k}z_{r}+e)$ and $h\in\left(L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\right)^{c}$. Hence ${\langle U_{g}w_{i},\sum_{r=1}^{k}z_{r}+e\rangle}\neq{\langle U_{h}w_{i},\sum_{r=1}^{k}z_{r}+e\rangle}$.
@@
[…]
@@ 2.2.2 Positivity of the Lower Lipschitz Constant
  • Next we construct inductively a sequence of non-zero vectors $z_{2},z_{3},...,z_{k}$ so that the local lower Lipschitz constant vanishes in a convex set of the form $\{\sum_{r=1}^{k}a_{r}z_{r}\leavevmode\nobreak\ ,\leavevmode\nobreak\ |a_{r}-1|<\delta\}$ for some $\delta>0$ small enough, and where sets $L^{i,j}$ remain constant. These steps are depicted in Figure 3 that suggests a "robotic arm" procedure (this name was suggested by the authors of [qaddura2024stable]).
@@
Applying Lemma 2.9 to two sequences $(x_{k})_{k}$ and $(y_{k})_{k}$ that satisfy (14-17) in 2.8, it follows that $d(x_{k},z_{1})=\lVert x_{k}-z_{1}\rVert$ and $d(y_{k},z_{1})=\lVert y_{k}-z_{1}\rVert$ for $k$ large enough. Hence alignment must occur from some rank on.
@@
For fixed $i\in[p]$, $j\in S_{i}$ and two sequences $(x_{n})_{n}$, $(y_{n})_{n}$ produced by 2.8, we denote by $g_{1,n,i,j}$ the group elements that achieves $\Phi_{i,j}(x_{n})$ and by $g_{2,n,i,j}$ the group element that achieves the $\Phi_{i,j}^{j}(y_{n})$. That is $\Phi_{i,j}(x_{n})={\langle U_{g_{1,n,i,j}}w_{i},x_{n}\rangle}$ and $\Phi_{i,j}(y_{n})={\langle U_{g_{2,n,i,j}}w_{i},y_{n}\rangle}$.
@@
For $i=1,\leavevmode\nobreak\ j=1$ there is a subsequence $(x_{n_{m}})_{m}$ such that $g_{1,1,1,n_{m}}=g_{1,1,1}$ for every $m\in\mathbb{N}$. Similarly, for $i=1,\leavevmode\nobreak\ j=2$ we can find a subsequence of $(x_{n_{m}})_{m}$, lets call it $(x_{n_{l}})_{l}$, such that $g_{1,1,2,n_{l}}=g_{1,1,2},\leavevmode\nobreak\ \forall l\in\mathbb{N}$. So by induction after $\sum_{i\in[p]}m_{i}=m$ steps we construct a subsequence of $(x_{n})_{n}$ lets call it $(x_{n_{m}})_{m}$ such that $g_{1,i,j,n_{m}}=g_{1,i,j}$ for every $i\in[p],j\in S_{i}$. Starting from sequence $(y_{1,n_{m}})_{m}$ we repeat the same procedure concluding in a subsequence $(y_{1,n_{r}})_{r}$ such that $g_{2,i,j,n_{r}}=g_{2,i,j}$ for every $r\in\mathbb{N},i\in[p],j\in S_{i}$ Notice that sequences $(x_{n_{r}})_{r}$ and $(y_{n_{r}})_{r}$ that from now on we will call them $(x_{n})_{n}$ and $(y_{n})_{n}$ for easiness of notation, satisfy the assumptions of lemma. ∎
@@
This sequence converge to 00, as $k\to\infty$ while also $u_{n},v_{n}\to 0$. So we conclude that for each $i\in[p]$ and $j\in S_{i}$, ${\langle U_{g_{1,i,j}}w_{i}-U_{g_{2,i,j}}w_{i},z_{1}\rangle}=0$. So
@@
where $\lVert u_{n}\rVert,\lVert v_{n}\rVert\to 0$, so for large enough $n$ we have that $\lVert u_{n}\rVert,\lVert v_{n}\rVert\leq\frac{1}{4}\rho_{0}(z_{1})$. Recall that from 2.9, we conclude that exists $N_{0}\in\mathbb{N}$, such that $\lVert u_{n}-v_{n}\rVert\leq\lVert U_{g}u_{n}-v_{n}\rVert$ for all $g\in H(z_{1})$ and $k\geq N_{0}$.
@@
    Thus from part (3) of 2.6, 2.3 and part (2) of 2.6 we have that for any $n\geq M_{0}$ and $(i,j)\in S$
@@
    Let two sequences $(x_{n})_{n},(y_{n})_{n}$ that both converge to $\sum_{r=1}^{k}z_{r}$, and achieve lower Lipschitz bound zero for map $\Phi_{{\textbf{w}},S}$. We align sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ to satisfy the properties of 2.8. We denote by $a_{n}=P_{E_{k}}x_{n}$ and $b_{n}=P_{E_{k}}y_{n}$ the orthogonal projections of the sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ respectively, on the linear subspace $E_{k}=\operatorname{span}\{z_{1},\dots,z_{k}\}^{\bot}$.
    Claim 1: First we will show that $\exists M_{0}$ such that $\forall n\geq M_{0}$, $a_{n}\neq 0$ or $b_{n}\neq 0$. Assuming otherwise, there are two sequences of vectors $x_{n}=\sum_{r=1}^{k}c_{r,n}z_{r}$ and $y_{n}=\sum_{r=1}^{k}d_{r,n}z_{r}$, where $\lim_{n\to\infty}c_{r,n}=\lim_{n\to\infty}d_{r,n}=1,\leavevmode\nobreak\ \forall r\in[k]$ that achieve lower Lipschitz bound zero. Recall that from part (2) of 2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
@@
    00 0 $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
@@
    Where the second equality comes from 2.6 part 3, third equality from 2.4 the fifth inclusion from 2.3, and the last equality from 2.6 part 2.
@@
    We prove this claim by contradiction. Assume that $\alpha\neq\delta$. From 2.9, $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$

### v4

# G-Invariant Representations using Coorbits: Bi-Lipschitz Properties

@@ 1 Introduction
In a lot of machine learning problems we want to embed our data into an Euclidean space $\mathbb{R}^{m}$ using a symmetry-invariant embedding $\Psi$ and utilize $\mathbb{R}^{m}$ as our feature space. This embedding $\Psi$ should also separate data orbits and satisfy certain stability conditions to ensure that small perturbations of the input do not significantly impact the predictions. We worked in the orbit separation problem in [11]. In this paper we focus on the stability problem.
This problem is an instance of invariant machine learning. [30, 5, 22, 16, 28, 20, 17, 19].
The most common group actions in invariant machine learning are permutations [33, 18, 10] reflections [29] and translations [25]. A related, by somewhat different problem is the case of equivariant embeddings [31, 28, 37, 36, 15].
The phase retrieval that was introduced in [9] is an instance of this setup, when the compact group is the torus. [8, 6, 7, 3, 24, 2].
Our work extends and unifies two prior approaches: the max filter embedding introduced in [18, 30], and the permutation invariant representation introduced in [10]. [4] analyzes the stability of generalized phase retrieval problems under the action of compact groups, providing conditions under which stable recovery is possible from invariant measurements. The construction of permutation invariant embeddings is closely connected to the phase retrieval problem [9, 1] that has a large body of results. For instance, [6] provides exact estimates for both the upper and lower Lipschitz bounds, addressing both the real and complex cases of the phase retrieval problem. In [13] we establish an isometric identification of the real phase retrieval problem to $S_{2}$-invariant representations.
A completely different approach is considered in [23]. There it is proved that for any discrete group $G$ of isometries acting on $\mathbb{R}^{d}$, it is possible to construct a bi-Lipschitz map from $\mathbb{R}^{d}/G$ to $\mathbb{R}^{N}$, where the distortion depends solely on $d$.
Since the first draft of this paper was placed on arxiv, the authors of [32] extended the construction of this paper and proved that for certain compact groups, given enough generic templates, the coorbit filter bank (as defined there) is injective and bi-Lipschitz.
@@ 1.1 Notation
In this paper we use the same notation as in [11]. Let $({\mathscr{V}},{\langle\cdot,\cdot\rangle})$ be a $d$-dimensional real vector space with scalar product, and $d\geq 2$. Assume $(G,\cdot)$ is a finite group of order $|G|=N$ acting unitarily on $\mathscr{V}$}. For every $g\in G$, we denote by $U_{g}x$ the group action on vector $x\in{\mathscr{V}}$. Let $\hat{{\mathscr{V}}}={\mathscr{V}}/\sim$ denote the quotient space with respect to this group action, $x\sim y$ if and only if $y=U_{g}x$ for some $g\in G$. We denote by $[x]$ the orbit of vector $x$, i.e. $x]=\{U_{g}x:g\in G\}$. The natural metric, $\operatorname{{\mathbf{d}}}:\hat{{\mathscr{V}}}\times\hat{{\mathscr{V}}}\to\mathbb{R}$, is defined by
@@
For an integer $p\in\mathbb{N}$, we denote $p]=\{1,2,\ldots,p\}$. For a set $S$, $|S|$ denotes its cardinal. Fix a $p$-tuple of vectors ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$. For any $i\in[p]$ and $j\in[N]$ we define the operator $\Phi_{w_{i},j}:{\mathscr{V}}\to\mathbb{R}$ so that $\Phi_{w_{i},j}(x)$ is the $j$-th coordinate of the sorted vector $\operatorname{\downarrow}({\langle U_{g}w_{i},x\rangle})_{g\in G}$. Fix a set $S\subset[N]\times[p]$ such that $|S|=m$, and for $i\in[p]$, let $S_{i}=\{k\in[N]:(k,i)\in S\}$ (the $i^{th}$ column of $S$). Denote by $m_{i}$ the cardinal of the set $S_{i}$, $m_{i}=|S_{i}|$. Thus $m=\sum_{i=1}^{p}m_{i}$.
@@
[…]
@@
Informally, our main results, ˜1.4, ˜1.6, ˜1.7 and ˜1.8, state that: (1) "injectivity" implies "(bi-Lipschitz) stability", (2) stable bi-Lipschitz embedding can be achieved into an Euclidean space of dimension at most twice the dimension of the input space; and (3) any continuous or even Lipschitz map factors through $\Phi_{{\textbf{w}},S}$.
@@ 1.2 Main results
[…]
Remark 1.9.
In [14] we generalize ˜1.4 for the case where our embedding has the form $a\circ\Phi_{{\textbf{w}},S}(x)$, where $a$ is a linear map.
Remark 1.10.
In [12] we study under which conditions the map $\hat{\Phi}_{{\textbf{w}},S}:\hat{{\mathscr{V}}}\to\mathbb{R}^{m}$ is injective on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/G$.
@@ 2 Stability of Embedding
Suppose that for ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$ the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective. In this case, we claim the map $\hat{\Phi}_{{\textbf{w}},S}$ is also bi-Lipschitz, (˜1.4). An estimate of the upper Lipschitz constant $b$ is given in ˜2.1.
˜1.5 follows from ˜1.4.
The proof of ˜1.4 is contained in the next two subsections.
@@ 2.1 Upper Lipschitz bound
$\lVert\operatorname{\downarrow}\{\Phi_{i,j}(x)\}_{j\in S_{i}}-\operatorname{\downarrow}\{\Phi_{i,j}(y)\}_{j\in S_{i}}\rVert$ $\leq\int_{0}^{1}(\sum_{j\in S_{i}}{\langle y-x,U_{g_{j_{t}}}w_{i}\rangle}^{2})^{1/2}dt$
@@ 2.2.1 Geometric Analysis of Coorbits
[…]
@@
then there exist at least Nj+1N-j+1, distinct elements of group $G$, say $\{h_{1},\dots h_{j}\}$ such that ${\langle U_{h_{k}}w_{i},x\rangle}<{\langle U_{g}w_{i},x\rangle}\leavevmode\nobreak\ \forall k\in[N-j+1]$, which is also a contradiction.
@@
But the set $\cup_{k\leq j}L^{i,k}(x)$ contains at least $j$ elements (since each $L^{i,j}(x)$ is non-empty) and so we derived a contradiction with ˜2.2(a) Equation˜6.
@@
Assume that exist $g_{1}\in L^{i,j}(x+c_{2}y)$ with $g_{1}\notin L^{i,j}(x+c_{1}y)$. Without loss of generality assume that ${\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}<\Phi_{i,j}(x+c_{1}y)$. Let $q>j$ be the smallest integer such that $g_{1}\in L^{i,q}(x+c_{1}y)$. Then $\Phi_{i,q}(x+c_{1}y)={\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}<\Phi_{i,j}(x+c_{1}y)$. By ˜2.2 (b)(ii),
@@
and $g_{1}\not\in\cup_{r\leq j}L^{i,r}(x+c_{1}y)$. On the other hand, from ˜2.2 (a), Equation˜6,
@@
1. Note that by norm conditions on $z_{r}$, ˜2.3 and ˜2.4, if $g_{1},g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$ then ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}\leavevmode\nobreak\ \forall l\in[k]$. Starting with $l=1$ and proceeding recursively we get ${\langle U_{g_{1}}w_{i},z_{r}\rangle}={\langle U_{g_{2}}w_{i},z_{r}\rangle}\leavevmode\nobreak\ \forall r\in[k]$. Therefore ${\langle U_{g_{1}}w_{i},a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},a_{r}z_{r}\rangle}\leavevmode\nobreak\ \forall r\in[k]$ and so ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}.$
@@
$\subset"$. From ˜2.3 we have that $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$.
$\supset"$. Let $g_{1}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$. Take any $g_{2}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. By ˜2.3, $g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$, but from part 1 of this Lemma,
@@
Case 1. Suppose $L^{i,j}(\sum_{r=1}^{k}z_{r})=G$. From ˜2.3 we know that
@@ 5. We prove the equality between complements:
First notice that by ˜2.3 we have that $L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$ and $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. From ˜2.6 part (2) we have that $L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. Take $g\in L^{i,j}(\sum_{r=1}^{k}z_{r}+e)$ and $h\in\left(L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\right)^{c}$. Hence ${\langle U_{g}w_{i},\sum_{r=1}^{k}z_{r}+e\rangle}\neq{\langle U_{h}w_{i},\sum_{r=1}^{k}z_{r}+e\rangle}$.
@@
$\geq\lVert w_{i}\rVert\left[\Delta^{i,j}(\sum_{r=1}^{k}z_{r})-2\left(\sum_{r=1}^{k}|a_{r}-1|\lVert z_{r}\rVert+\lVert e\rVert\right)\right]>0\hskip 56.9055pt(*)$
@@ 2.2.2 Positivity of the Lower Lipschitz Constant
  • Next we construct inductively a sequence of non-zero vectors $z_{2},z_{3},...,z_{k}$ so that the local lower Lipschitz constant vanishes in a convex set of the form $\{\sum_{r=1}^{k}a_{r}z_{r}\leavevmode\nobreak\ ,\leavevmode\nobreak\ |a_{r}-1|<\delta\}$ for some $\delta>0$ small enough, and where sets $L^{i,j}$ remain constant. These steps are depicted in Figure 3 that suggests a "robotic arm" procedure (this name was suggested by the authors of [32]).
@@
Applying Lemma 2.9 to two sequences $(x_{k})_{k}$ and $(y_{k})_{k}$ that satisfy (14-17) in ˜2.8, it follows that $d(x_{k},z_{1})=\lVert x_{k}-z_{1}\rVert$ and $d(y_{k},z_{1})=\lVert y_{k}-z_{1}\rVert$ for $k$ large enough. Hence alignment must occur from some rank on.
@@
For fixed $i\in[p]$, $j\in S_{i}$ and two sequences $(x_{n})_{n}$, $(y_{n})_{n}$ produced by ˜2.8, we denote by $g_{1,n,i,j}$ the group elements that achieves $\Phi_{i,j}(x_{n})$ and by $g_{2,n,i,j}$ the group element that achieves the $\Phi_{i,j}^{j}(y_{n})$. That is $\Phi_{i,j}(x_{n})={\langle U_{g_{1,n,i,j}}w_{i},x_{n}\rangle}$ and $\Phi_{i,j}(y_{n})={\langle U_{g_{2,n,i,j}}w_{i},y_{n}\rangle}$.
@@
For $i=1,\leavevmode\nobreak\ j=1$ there is a subsequence $(x_{n_{m}})_{m}$ such that $g_{1,1,1,n_{m}}=g_{1,1,1}$ for every $m\in\mathbb{N}$. Similarly, for $i=1,\leavevmode\nobreak\ j=2$ we can find a subsequence of $(x_{n_{m}})_{m}$, lets call it $(x_{n_{l}})_{l}$, such that $g_{1,1,2,n_{l}}=g_{1,1,2},\leavevmode\nobreak\ \forall l\in\mathbb{N}$. So by induction after $\sum_{i\in[p]}m_{i}=m$ steps we construct a subsequence of $(x_{n})_{n}$ lets call it $(x_{n_{m}})_{m}$ such that $g_{1,i,j,n_{m}}=g_{1,i,j}$ for every $i\in[p],j\in S_{i}$. Starting from sequence $(y_{1,n_{m}})_{m}$ we repeat the same procedure concluding in a subsequence $(y_{1,n_{r}})_{r}$ such that $g_{2,i,j,n_{r}}=g_{2,i,j}$ for every $r\in\mathbb{N},i\in[p],j\in S_{i}$ . Notice that sequences $(x_{n_{r}})_{r}$ and $(y_{n_{r}})_{r}$ that from now on we will call them $(x_{n})_{n}$ and $(y_{n})_{n}$ for easiness of notation, satisfy the assumptions of lemma. ∎
@@
This sequence converge to 0, as $k\to\infty$ while also $u_{n},v_{n}\to 0$. So we conclude that for each $i\in[p]$ and $j\in S_{i}$, ${\langle U_{g_{1,i,j}}w_{i}-U_{g_{2,i,j}}w_{i},z_{1}\rangle}=0$. So
@@
where $\lVert u_{n}\rVert,\lVert v_{n}\rVert\to 0$, so for large enough $n$ we have that $\lVert u_{n}\rVert,\lVert v_{n}\rVert\leq\frac{1}{4}\rho_{0}(z_{1})$. Recall that from ˜2.9, we conclude that exists $N_{0}\in\mathbb{N}$, such that $\lVert u_{n}-v_{n}\rVert\leq\lVert U_{g}u_{n}-v_{n}\rVert$ for all $g\in H(z_{1})$ and $k\geq N_{0}$.
@@
    Thus from part (3) of ˜2.6, ˜2.3 and part (2) of ˜2.6 we have that for any $n\geq M_{0}$ and $(i,j)\in S$
@@
    Let two sequences $(x_{n})_{n},(y_{n})_{n}$ that both converge to $\sum_{r=1}^{k}z_{r}$, and achieve lower Lipschitz bound zero for map $\Phi_{{\textbf{w}},S}$. We align sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ to satisfy the properties of ˜2.8. We denote by $a_{n}=P_{E_{k}}x_{n}$ and $b_{n}=P_{E_{k}}y_{n}$ the orthogonal projections of the sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ respectively, on the linear subspace $E_{k}=\operatorname{span}\{z_{1},\dots,z_{k}\}^{\bot}$.
    Claim 1: First we will show that $\exists M_{0}$ such that $\forall n\geq M_{0}$, $a_{n}\neq 0$ or $b_{n}\neq 0$. Assuming otherwise, there are two sequences of vectors $x_{n}=\sum_{r=1}^{k}c_{r,n}z_{r}$ and $y_{n}=\sum_{r=1}^{k}d_{r,n}z_{r}$, where $\lim_{n\to\infty}c_{r,n}=\lim_{n\to\infty}d_{r,n}=1,\leavevmode\nobreak\ \forall r\in[k]$ that achieve lower Lipschitz bound zero. Recall that from part (2) of ˜2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
@@
    $0$ $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
@@
    Where the second equality comes from ˜2.6 part 3, third equality from ˜2.4 the fifth inclusion from ˜2.3, and the last equality from ˜2.6 part 2.
@@
    We prove this claim by contradiction. Assume that $\alpha\neq\delta$. From ˜2.9, $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$